The number of short-term parking spaces at 13 airports is shown. Round your answers to one decimal place. 1882 7807 6469 1754 8854 5270 9210 1190 2434 6273 7546 2601 8890 Source: USA TODAY Download data Part 1 out of 4 Find the mean. LOADING 0% Mean =
Answer #1

\large \mathbf{Solution}

X
1882
7807
6469
1754
8854
5270
9210
1190
2434
6273
7546
2601
8890
\sum X=70180

\mathbf{Mean:}\frac{\sum X}{n}=\frac{70180}{13}=\mathbf{{\color{Red} 5398.5}}